A firm has a positive economic value added (EVA) of $500 (all amounts in thousands), long-term...

  1. A firm has a positive economic value added (EVA) of $500 (all amounts in thousands), long-term debt of $20,000, equity of $40,000, a weighted average cost of capital (WACC) of 9% and a marginal tax rate of 32%. What is this organization's earnings before interest and taxes (EBIT)?

  1. $4,012
  2. $5,900
  3. $8,676
  4. $18,438
Answer #1

Economic value added = $500

Long term debt = $20,000

Equity = $40,000

WACC = 9%

Marginal tax rate = 32%

EBIT = ?

Capital employed = Long term debt + Equity

= 20,000 + 40,000

= $60,000

Economic value added = Operating profit after tax - Cost of capital

500 = Operating profit after tax - 60,000 x 9%

500 = Operating profit after tax - 5,400

Operating profit after tax = $5,900

EBIT = Operating profit after tax/(1 - Tax rate)

= 5,900/(1- 0.32)

= 5,900/0.68

= $8,676

Correct option is (C)

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.